OpenStudy (anonymous):

PLEASE HELP! How many liters of a 2.25 molar hydrobromic acid (HBr) solution would be needed to react completely with 100.0 grams of calcium metal? Ca (s) + 2HBr (aq) CaBr2 (aq) + H2 (g)

OpenStudy (anonymous):

I got 2.00? the answer choices are 1.11 liters 2.22 liters 2.81 liters 22.5 liters

OpenStudy (aaronq):

convert 100 g of calcium to moles then use the concentration of the HBr in the molarity equation (don't forget to take the stoichiometric coefficient into account)

OpenStudy (anonymous):

so it would be: 100g Ca x 1 mol Ca/ 40.1 g Ca x 2 mol HBr/1 mol Ca x 1 L HCl solution/ 2.5 mol HBr

OpenStudy (aaronq):

nope, do one step at a time.. molarity of Hbr = the moles of calciumx2/L you're solving for liters

OpenStudy (aaronq):

wait you're actually right

OpenStudy (anonymous):

but my answer is 2.00???

OpenStudy (aaronq):

but it's 2.25 M not 2.5

OpenStudy (anonymous):

oh...wow:/

OpenStudy (aaronq):

lol

OpenStudy (anonymous):

will you help me with one more?

OpenStudy (aaronq):

sure

OpenStudy (anonymous):

The density of an unknown gas is found to be 3.44 g/L at standard temperature and pressure (STP). What is the molar mass of this gas?

OpenStudy (aaronq):

use the ideal gas law and manipulate it to solve for molar mass

OpenStudy (anonymous):

how do I do that..?

OpenStudy (aaronq):

algebra and substitution

OpenStudy (aaronq):

try it, if you can't do it ill show you

OpenStudy (anonymous):

well idk how to start it

OpenStudy (anonymous):

ik ideal gas law is Pv=nRT

OpenStudy (aaronq):

PV=nRT moles definition: n=m/M PV=mRT/M

OpenStudy (anonymous):

3.44g/L x 0.0821L-atm/mole-K x 273K / 1atm = molar mass = 77.1g/mole?

OpenStudy (aaronq):

looks right

OpenStudy (anonymous):

ok thank you!!

OpenStudy (aaronq):

no prob!